OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = 3x + 2, what is f(g(x)) ?

OpenStudy (shubhamsrg):

f(g(x)) means f(3x+2) which is 4(3x+2) +1

OpenStudy (lgbasallote):

like in f(6) you replace the x of f(x) with 6 right? so in f(g(x)) you replace the x of f(x) with g(x) so it becomes 4(g(x)) + 1

OpenStudy (anonymous):

f0g=4(3x+2)+1 fog=12x+8+1 =12x+9

OpenStudy (anonymous):

replace any x in the f(x)=(4x+1) with g(x)=(3x+2) f(g(x)) = 4(3x+2)+1 = 12x+8+1 = 12x+9 = 3(4x+3)

OpenStudy (anonymous):

yeah............Zara26